Prize Mustard uses a process costing system with two departments: (a) a Mixing Department, and (b) a Bottling Department. During May, Prize generated $120,000 in revenue by producing 40,000 gallons (10,000 cases) of mustard. The companyreduces its work in process inventories to zero eachmonth. Manufacturing costs for May are as follows Mixing Bottling Department Department Direct materials Direct labor Manufacturing overhea $16,000 $5,600 $11,400 $6,000 S3,900 $8,100 1 Refer to the above data. The unit cost per gallon of mustard processed by the Mixing Department in May was $ 2 Refer to the above data. The unit cost per case of mustard incurred by the Bottling Department in May was $ Refer to the above data. The total cost ofthe 10,000 cases transferred to finished goods in May was S 3 4 Refer to the above data. Assume that 9,500 cases of mustard were sold in May The amount charged to cost of goods sold in May S 5 Refer to the above data. Assume that $9,000 of finished goods were on hand at May 1. Under the LIFO method of inventory valuation, finished goods inventory at May 31 was S if 9.500 cases of mustard were sold in May
Answer #1
1

total cost in the mixing department = $33000

unit cost per gallon = 33000/40000 =$0.825

16000+15600+11400
2

total cost in the bottling department = $18000

unit cost per case = 18000/10000 = $1.8

6000+3900+8100
3

total cost of production in both department = $51000

33000(16000+15600+11400)+

18000(6000+3900+8100)

4

if 9500 units were sold the cost of goods charged will be

(total cost/total number of units produced) *no. of units sold

= 51000/10000*9500 = $48450

only the part which is actually sold is transferred from the cost of goods sold account
5

if 9500 are sold off and there was an opening stock o 9000 and the company followed LIO method then

the balance in the finished goods account will be 9500 units  

total cost/number of units made *no. of units of closing stock = 51000/10000*9500 =48450

it is assumed that the cost was the same in the last month as there is no information given